OpenStudy (ronlover101):

Hey can someone help me with this tricky math equation? I'm a freshman doing Algebra 1.

OpenStudy (ronlover101):

(x+3) +1 = 2

OpenStudy (will.h):

Solving for x?

OpenStudy (ronlover101):

yes

OpenStudy (will.h):

Okay We will have to simplify the left side We can get rid of the parentheses since the parentheses are not being multiplied by another term So we have X + 3 + 1 = 2

OpenStudy (ronlover101):

I have the answers from a test guide that our teacher gave us... But I don't know how to get there. Here is the two answers. x=-2 and x=-4

OpenStudy (sooobored):

are you sure its a parenthesis () and not absolute values | |

OpenStudy (shadowlegendx):

\[(x+3) +1 = 2\] We can remove the parentheses as @Will.H stated. \[x + 3 + 1 = 2\] Like terms, \[x + 4 = 2\] Subtract 4 from both sides \[x = -2\]

OpenStudy (shadowlegendx):

@sooobored good point

OpenStudy (ronlover101):

the are like two brackets

OpenStudy (ronlover101):

but I can't type them

OpenStudy (will.h):

Absolute values then

OpenStudy (shadowlegendx):

\[\left| x + 3 \right| + 1 = 2\]

OpenStudy (will.h):

Since there are 2 solutions

OpenStudy (shadowlegendx):

Like this right?

OpenStudy (ronlover101):

Ya you were right on the first one @ShadowLegendX

OpenStudy (ronlover101):

wait

OpenStudy (ronlover101):

we were told to do the opposite if it was adding we had to subtract. so wouldn't we need to subtract the 1?

OpenStudy (shadowlegendx):

Lets solve \[\left| x + 3 \right| + 1 = 2 \] Subtract 1 from both sides to isolate x \[\left| x + 3 \right| = 1\] Now we get two equations \[x + 3 = 1\] and \[x + 3 = -1\] Solving for the first one \[x + 3 = 1\] Subtract 3 from both sides \[x = -2\] Solving for the second one \[x + 3 = -1\] Subtract 3 from both sides \[x = -4\]

OpenStudy (will.h):

@ShadowLegendX brilliant work mate

OpenStudy (ronlover101):

thank you so much!!

OpenStudy (shadowlegendx):

You can tell that these are solutions to, \[\left| x + 3 \right| + 1 = 2\] Because if you substitute either of the solutions for x, and solve, you get 2.

OpenStudy (ronlover101):

now i remember how to do this!!! You are a big help thank you!!

OpenStudy (shadowlegendx):

No problem, glad I could help :)